# XLSR-MamBo: Scaling the Hybrid Mamba-Attention Backbone for Audio Deepfake Detection

### v1

@@ 1 Introduction
Driven by rapid advances in generative artificial intelligence (GenAI), cutting-edge text-to-speech (TTS) and voice conversion (VC) technologies have achieved remarkable progress (casanova2024xtts; du2025cosyvoice; zhou2025indextts2). The naturalness and timbre similarity of synthesized speech have reached levels that are increasingly challenging for the human ear to distinguish. While beneficial for applications like audiobooks and content creation, these technologies pose severe risks if misused (li2025survey). The maturation of zero-shot and few-shot voice cloning allows attackers to synthesize highly realistic forged speech using only short audio segments collected from social media (azzuni2025voice). Such misuse not only facilitates the spread of misinformation and erodes societal trust but also threatens voice biometric systems (barrington2025people; DeepfakesEUNews). Consequently, developing robust and effective audio deepfake detection (ADD) models is imperative.
@@
Early research in anti-spoofing primarily relied on handcrafted features to detect synthesis artifacts that deviate from human speech production mechanisms (khan2023battling; yang2025forensic). Although such methods offer interpretability and computational efficiency, these often lack generalization against unseen advanced generators. With the advancement of deep learning, mainstream ADD models have shifted to end-to-end architectures, which directly process raw audio for feature extraction and classification. Regarding front-end feature extraction, XLSR (babu2022xlsr) is a self-supervised learning (SSL) model based on Wav2Vec 2.0 (baevski2020wav2vec), pre-trained on massive cross-lingual datasets. Benefiting from this large-scale diversity, XLSR has demonstrated competitive generalization capabilities compared to other pre-trained models (PTMs) such as Whisper (radford2023robust) and WavLM (chen2022wavlm), demonstrating strong competitiveness in ADD tasks (kashyap2025fooling). Building on this powerful representation, attention-based classifiers have become an efficient back-end candidate. The Conformer architecture (gulati2020conformer), which combines the global context modeling of Transformers with the local feature capture of CNNs, is widely adopted. By integrating XLSR with Conformer, the XLSR-Conformer (rosello2023conformer) achieved state-of-the-art (SOTA) performance on the ASVspoof 2021 LA (ASV21LA) and DF (ASV21DF) (yamagishi2021asvspoof) dataset at the time. Subsequent variants, such as XLSR-Conformer+TCM truong2024temporal, XLSR-SLS zhang2024audio, and Kanformer (dat2025xlsr), further pushed the performance boundaries by refining the attention mechanism.
@@
However, the quadratic computational complexity of the classical Transformer limits its efficiency in processing long audio sequences. To address this, state space models (SSMs) (gu2022efficiently) with linear complexity, particularly the Mamba architecture (gu2023mamba), have attracted widespread attention. RawBMamba (chen2024rawbmamba) was among the early works to explore the application of Mamba in audio deepfake detection. Addressing the unidirectional limitations of the original Mamba, it proposed a Bidirectional Mamba architecture that combines SincLayers/convolutional layers (to capture short-term local features) with bidirectional Mamba blocks (to capture long-range contextual dependencies). This design effectively identifies subtle artifacts in synthesized audio and significantly outperforms traditional end-to-end models like Rawformer (liu2023leveraging), validating the generalization capability of the Mamba architecture in audio forensics. Following this, works like XLSR-Mamba (xiao2025xlsr) and Fake-Mamba (xuan2025fake) further advanced the field. However, a critical observation in these works is their reliance on manually designed bidirectional strategies (e.g., dual-branch fusion) to approximate the non-causal receptivity of Transformers. While these studies successfully validated that substituting Transformer layers with SSM blocks yields efficiency gains, relying solely on this replacement strategy overlooks the unique complementary strengths of different mechanisms.
@@
Although pure SSM models excel in inference efficiency and ADD performance, experimental results from cutting-edge hybrid architectures such as Zamba (glorioso2024zamba), Samba (ren2024samba), Jamba (lieber2024jamba), and Nemotron-H (blakeman2025nemotron) demonstrate that combining SSMs with Attention yields competitive performance. Furthermore, hardware-aware algorithms such as FlashAttention (dao2022flashattention) and Mesh-Attention (chen2025mesh) have alleviated the computational bottlenecks of Transformers. From the perspective of inductive bias, SSMs specialize in information compression and temporal recurrence, whereas Attention excels at content-based, precise retrieval and associative recall via the induction heads mechanism (olsson2022context; arora2024simple). Consequently, in ADD tasks, the Attention mechanism can be better suited to globally correlate specific frequency-domain forgery artifacts. Specifically, deepfake signals often manifest as a duality: subtle local high-frequency artifacts (requiring the fine-grained temporal recurrence of SSMs) and global spectral inconsistencies (requiring the global retrieval of Attention). Relying solely on one mechanism limits the model’s ability to capture this full spectrum of manipulation traces (frank2020leveraging; sun2023ai).
@@
In light of this, we propose a novel framework, XLSR-MamBo, for anti-spoofing. This framework employs XLSR as the front-end and utilizes a hybrid SSM-Attention architecture as the back-end encoder. Compared to prior works that rely on manual bidirectionality, we explore advanced variants including Mamba2 (dao2024transformers), Gated DeltaNet (GDN) (yang2024gated), and notably, Hydra (hwang2024hydra). Hydra introduces a paradigm shift by parameterizing the sequence mixer as a quasiseparable matrix. Unlike heuristic dual-branch extensions, this formulation allows Hydra to perform native bidirectional processing with linear complexity, capturing holistic non-causal dependencies without structural redundancy. This theoretical advantage positions it as an ideal candidate for detecting artifact traces that violate causal consistency. Furthermore, we investigate the impact of scaling the backbone depth to mitigate inference instability. The proposed framework is evaluated on the ASV21LA and DF, In-the-Wild (ITW) (muller2022does), and DFADD (du2024dfadd) datasets, and the results demonstrate its robustness and effectiveness compared to other SOTA systems.
The code and trained models will be released to the community once finalized.
@@ 2 Methods
The hybrid design is motivated by the complementary strengths of SSMs (efficient sequence modeling) and Attention mechanisms (precise global context recall). Inspired by MambaFormer (park2024can), we place the Mamba block after the input projection layer. This configuration leverages the implicit positional encoding inherent in the recurrent structure of SSMs, mitigating the need for explicit positional encodings while maintaining robust sequence modeling (gu2023mamba). All blocks incorporate Pre-Norm (xiong2020layer) connections to enhance training stability.
@@
MamBo-1 (Mamba Layer, Figure 1 (a)): Replaces the multi-head self-attention (MHA) with a Mamba block while retaining the subsequent SwiGLU shazeer2020glu feed-forward network (FFN). This provides a baseline to isolate the standalone effectiveness of SSM variants in capturing long-range dependencies without explicit attention.
@@
MamBo-3 (Mamba-Transformer Layer, Figure 1 (c)): Adopts a classic interleaved strategy by alternating standard Mamba Layers and Transformer Layers. Building on recent works (zhang2025mamba) that position Mamba as a promising alternative to self-attention in speech modeling. This sequential hybridization provides a baseline for evaluating the benefits of separating SSM processing and Attention-based refinement into distinct layers.
@@
As shown in Figure 1, the overall pipeline of the proposed XLSR-MamBo is structured as follows. The input raw waveform is first processed by the pre-trained XLSR model, which extracts a sequence of high-level speech representations, resulting in a feature matrix $X\in\mathbb{R}^{T\times 1024}$, where $T$ denotes the number of time frames proportional to the input audio duration. To stabilize training and align feature dimensionality, we apply RMSNorm (zhang2019root) followed by a linear projection layer, mapping the channel dimension from 1024 to the hidden dimension $D$. The resulting projected features $H\in\mathbb{R}^{T\times D}$ serve as input to the backbone encoder, which consists of stacked MamBo layers ($L$). This encoder effectively captures complex temporal dependencies and subtle manipulation artifacts while preserving the feature shape $(T\times D)$. Subsequently, the frame-level features are aggregated into a fixed-dimensional utterance-level representation using gated attention pooling. Finally, a single linear layer projects this pooled representation to a 2-dimensional output, yielding logits for binary classification (spoof vs. bonafide).
@@ 3.1 Datasets and Metrics
All MamBo variants were trained on the ASVspoof 2019 LA (ASV19LA) training set (wang2020asvspoof), which comprises 25,380 training, 24,844 development, and 71,237 evaluation utterances. To evaluate the generalizability and robustness of the proposed method via cross-dataset evaluation, we conducted experiments on four challenging test sets: ASV21LA (181,566 utterances), ASV21DF (611,829 utterances), ITW (31,778 utterances), and DFADD (1,355 utterances each subset). The ASVspoof 2021 introduces complex transmission channel variations and compression artifacts. The ITW dataset contains real-world deepfake audio sourced from social media, while DFADD includes synthetic speech generated by advanced diffusion- and flow-matching-based TTS models (subsets D1–D3 and F1–F2, respectively). Following the standard protocols of the ASVspoof Challenges, we report the equal error rate (EER) and the minimum tandem detection cost function (min t-DCF) as evaluation metrics.
@@ 3.2 Implementation Details
For data augmentation, we used RawBoost (tak2022rawboost). Consistent with baseline configurations (xiao2025xlsr; xuan2025fake), we trained separate models for the ASV21LA (21LA) and ASV21DF (21DF) evaluations. The 21LA model applied linear and non-linear convolutive noise combined with impulsive signal-dependent noise, while the 21DF model used stationary signal-independent noise.
All audio inputs were processed at 16 kHz with a fixed duration of 4.175s (66,800 samples). We fine-tuned the models using the AdamW (loshchilov2017decoupled) optimizer ($lr=10^{-5}$, $\beta_{1}=0.9,\beta_{2}=0.95$, weight decay=0.05) with a scheduler combining 10% linear warmup with cosine decay. Training was performed in mixed precision (BF16/FP32) using FocalLoss (lin2017focal) to mitigate class imbalance, for a maximum of 20 epochs with early stopping (patience of 7 epochs), and a batch size of 32. The top-5 checkpoints with the lowest validation loss were retained. For the architecture, the projection dimension was set to $D=128$, with $L=5$ MamBo layers by default. For the Mamba variants (Mamba2, Hydra, GDN), the state dimension (d_state) and head dimension were set to 64 and 32, respectively. All experiments were conducted on one NVIDIA RTX 4080 Super 16 GB (training approximately 1.5 hours per model) and two NVIDIA RTX 3090 24 GB GPUs. To facilitate reproducibility, a fixed random seed was used in experimental runs.
@@ 4.3 Compare with Other SOTA Systems
    chen2024rawbmamba; 2222kheir2025bicrossmamba; 3333rosello2023conformer; 4444truong2024temporal; 5555zhang2024audio; 6666xiao2025xlsr; 7777xuan2025fake.

### v2

@@ 1 Introduction
Driven by rapid advances in generative artificial intelligence (GenAI), cutting-edge text-to-speech (TTS) and voice conversion (VC) technologies have achieved remarkable progress (Casanova et al., 2024; Du et al., 2025; Zhou et al., 2025). The naturalness and timbre similarity of synthesized speech have reached levels that are increasingly challenging for the human ear to distinguish. While beneficial for applications like audiobooks and content creation, these technologies pose severe risks if misused (Li et al., 2025). The maturation of zero-shot and few-shot voice cloning allows attackers to synthesize highly realistic forged speech using only short audio segments collected from social media (Azzuni and Saddik, 2025). Such misuse not only facilitates the spread of misinformation and erodes societal trust but also threatens voice biometric systems (Barrington et al., 2025; McConvey, 2025). Consequently, developing robust and effective audio deepfake detection (ADD) models is imperative.
@@
Early research in anti-spoofing primarily relied on handcrafted features to detect synthesis artifacts that deviate from human speech production mechanisms (Khan et al., 2023; Yang et al., 2026). Although such methods offer interpretability and computational efficiency, these often lack generalization against unseen advanced generators. With the advancement of deep learning, mainstream ADD models have shifted to end-to-end architectures, which directly process raw audio for feature extraction and classification. Regarding front-end feature extraction, XLSR (Babu et al., 2022) is a self-supervised learning (SSL) model based on Wav2Vec 2.0 (Baevski et al., 2020), pre-trained on massive cross-lingual datasets. Benefiting from this large-scale diversity, XLSR has demonstrated competitive generalization capabilities compared to other pre-trained models (PTMs) such as Whisper (Radford et al., 2023) and WavLM (Chen et al., 2022), demonstrating strong competitiveness in ADD tasks (Kashyap et al., 2025). Building on this powerful representation, attention-based classifiers have become an efficient back-end candidate. The Conformer architecture (Gulati et al., 2020), which combines the global context modeling of Transformers with the local feature capture of CNNs, is widely adopted. By integrating XLSR with Conformer, the XLSR-Conformer (Rosello et al., 2023) achieved state-of-the-art (SOTA) performance on the ASVspoof 2021 LA (ASV21LA) and DF (ASV21DF) (Yamagishi et al., 2021) dataset at the time. Subsequent variants, such as XLSR-Conformer+TCM Truong et al. (2024), XLSR-SLS Zhang et al. (2024), and Kanformer (Dat and Dat, 2025), further pushed the performance boundaries by refining the attention mechanism.
@@
However, the quadratic computational complexity of the classical Transformer limits its efficiency in processing long audio sequences. To address this, state space models (SSMs) (Gu et al., 2022) with linear complexity, particularly the Mamba architecture (Gu and Dao, 2024), have attracted widespread attention. RawBMamba (Chen et al., 2024) was among the early works to explore the application of Mamba in audio deepfake detection. Addressing the unidirectional limitations of the original Mamba, it proposed a Bidirectional Mamba architecture that combines SincLayers/convolutional layers (to capture short-term local features) with bidirectional Mamba blocks (to capture long-range contextual dependencies). This design effectively identifies subtle artifacts in synthesized audio and significantly outperforms traditional end-to-end models like Rawformer (Liu et al., 2023), validating the generalization capability of the Mamba architecture in audio forensics. Following this, works like XLSR-Mamba (Xiao and Das, 2025) and Fake-Mamba (Xuan et al., 2025) further advanced the field. However, a critical observation in these works is their reliance on manually designed bidirectional strategies (e.g., dual-branch fusion) to approximate the non-causal receptivity of Transformers. While these studies successfully validated that substituting Transformer layers with SSM blocks yields efficiency gains, relying solely on this replacement strategy overlooks the unique complementary strengths of different mechanisms.
@@
Although pure SSM models excel in inference efficiency and ADD performance, experimental results from cutting-edge hybrid architectures such as Zamba (Glorioso et al., 2024), Samba (Ren et al., 2025), Jamba (Lieber et al., 2025), and Nemotron-H (Blakeman et al., 2025) demonstrate that combining SSMs with Attention yields competitive performance. Furthermore, hardware-aware algorithms such as FlashAttention (Dao et al., 2022) and Mesh-Attention (Chen et al., 2025) have alleviated the computational bottlenecks of Transformers. From the perspective of inductive bias, SSMs specialize in information compression and temporal recurrence, whereas Attention excels at content-based, precise retrieval and associative recall via the induction heads mechanism (Olsson et al., 2022; Arora et al., 2024). Consequently, in ADD tasks, the Attention mechanism can be better suited to globally correlate specific frequency-domain forgery artifacts. Specifically, deepfake signals often manifest as a duality: subtle local high-frequency artifacts (requiring the fine-grained temporal recurrence of SSMs) and global spectral inconsistencies (requiring the global retrieval of Attention). Relying solely on one mechanism limits the model’s ability to capture this full spectrum of manipulation traces (Frank et al., 2020; Sun et al., 2023).
@@
In light of this, we propose a novel framework, XLSR-MamBo, for anti-spoofing. This framework employs XLSR as the front-end and utilizes a hybrid SSM-Attention architecture as the back-end encoder. Compared to prior works that rely on manual bidirectionality, we explore advanced variants including Mamba2 (Dao and Gu, 2024), Gated DeltaNet (GDN) (Yang et al., 2024), and notably, Hydra (Hwang et al., 2024). Hydra introduces a paradigm shift by parameterizing the sequence mixer as a quasiseparable matrix. Unlike heuristic dual-branch extensions, this formulation allows Hydra to perform native bidirectional processing with linear complexity, capturing holistic non-causal dependencies without structural redundancy. This theoretical advantage positions it as an ideal candidate for detecting artifact traces that violate causal consistency. Furthermore, we investigate the impact of scaling the backbone depth to mitigate inference instability. The proposed framework is evaluated on the ASV21LA and DF, In-the-Wild (ITW) (Müller et al., 2022), and DFADD (Du et al., 2024) datasets, and the results demonstrate its robustness and effectiveness compared to other SOTA systems.
@@ 2 Methods
The hybrid design is motivated by the complementary strengths of SSMs (efficient sequence modeling) and Attention mechanisms (precise global context recall). Inspired by MambaFormer (Park et al., 2024), we place the Mamba block after the input projection layer. This configuration leverages the implicit positional encoding inherent in the recurrent structure of SSMs, mitigating the need for explicit positional encodings while maintaining robust sequence modeling (Gu and Dao, 2024). All blocks incorporate Pre-Norm (Xiong et al., 2020) connections to enhance training stability.
@@
MamBo-1 (Mamba Layer, Figure 1 (a)): Replaces the multi-head self-attention (MHA) with a Mamba block while retaining the subsequent SwiGLU Shazeer (2020) feed-forward network (FFN). This provides a baseline to isolate the standalone effectiveness of SSM variants in capturing long-range dependencies without explicit attention.
@@
MamBo-3 (Mamba-Transformer Layer, Figure 1 (c)): Adopts a classic interleaved strategy by alternating standard Mamba Layers and Transformer Layers. Building on recent works (Zhang et al., 2025) that position Mamba as a promising alternative to self-attention in speech modeling. This sequential hybridization provides a baseline for evaluating the benefits of separating SSM processing and Attention-based refinement into distinct layers.
@@
As shown in Figure 1, the overall pipeline of the proposed XLSR-MamBo is structured as follows. The input raw waveform is first processed by the pre-trained XLSR model, which extracts a sequence of high-level speech representations, resulting in a feature matrix $X\in\mathbb{R}^{T\times 1024}$, where $T$ denotes the number of time frames proportional to the input audio duration. To stabilize training and align feature dimensionality, we apply RMSNorm (Zhang and Sennrich, 2019) followed by a linear projection layer, mapping the channel dimension from 1024 to the hidden dimension $D$. The resulting projected features $H\in\mathbb{R}^{T\times D}$ serve as input to the backbone encoder, which consists of stacked MamBo layers ($L$). This encoder effectively captures complex temporal dependencies and subtle manipulation artifacts while preserving the feature shape $(T\times D)$. Subsequently, the frame-level features are aggregated into a fixed-dimensional utterance-level representation using gated attention pooling. Finally, a single linear layer projects this pooled representation to a 2-dimensional output, yielding logits for binary classification (spoof vs. bonafide).
@@ 3.1 Datasets and Metrics
All MamBo variants were trained on the ASVspoof 2019 LA (ASV19LA) training set (Wang et al., 2020), which comprises 25,380 training, 24,844 development, and 71,237 evaluation utterances. To evaluate the generalizability and robustness of the proposed method via cross-dataset evaluation, we conducted experiments on four challenging test sets: ASV21LA (181,566 utterances), ASV21DF (611,829 utterances), ITW (31,778 utterances), and DFADD (1,355 utterances each subset). The ASVspoof 2021 introduces complex transmission channel variations and compression artifacts. The ITW dataset contains real-world deepfake audio sourced from social media, while DFADD includes synthetic speech generated by advanced diffusion- and flow-matching-based TTS models (subsets D1–D3 and F1–F2, respectively). Following the standard protocols of the ASVspoof Challenges, we report the equal error rate (EER) and the minimum tandem detection cost function (min t-DCF) as evaluation metrics.
@@ 3.2 Implementation Details
For data augmentation, we used RawBoost (Tak et al., 2022). Consistent with baseline configurations (Xiao and Das, 2025; Xuan et al., 2025), we trained separate models for the ASV21LA (21LA) and ASV21DF (21DF) evaluations. The 21LA model applied linear and non-linear convolutive noise combined with impulsive signal-dependent noise, while the 21DF model used stationary signal-independent noise.
All audio inputs were processed at 16 kHz with a fixed duration of 4.175s (66,800 samples). We fine-tuned the models using the AdamW (Loshchilov and Hutter, 2019) optimizer ($lr=10^{-5}$, $\beta_{1}=0.9,\beta_{2}=0.95$, weight decay=0.05) with a scheduler combining 10% linear warmup with cosine decay. Training was performed in mixed precision (BF16/FP32) using FocalLoss (Lin et al., 2017) to mitigate class imbalance, for a maximum of 20 epochs with early stopping (patience of 7 epochs), and a batch size of 32. The top-5 checkpoints with the lowest validation loss were retained. For the architecture, the projection dimension was set to $D=128$, with $L=5$ MamBo layers by default. For the Mamba variants (Mamba2, Hydra, GDN), the state dimension (d_state) and head dimension were set to 64 and 32, respectively. All experiments were conducted on one NVIDIA RTX 4080 Super 16 GB (training approximately 1.5 hours per model) and two NVIDIA RTX 3090 24 GB GPUs. To facilitate reproducibility, a fixed random seed was used in experimental runs.
@@ 4.3 Compare with Other SOTA Systems
    Chen et al. (2024); 2222Kheir et al. (2025); 3333Rosello et al. (2023); 4444Truong et al. (2024); 5555Zhang et al. (2024); 6666Xiao and Das (2025); 7777Xuan et al. (2025).